what is the function rule if:
x y
-2 = -2
0 = 1
y=???

Answer by stanbon(75887)

x y
-2 = -2
0 = 1
-----------------
If I read you right you have two points: (-2,-2) and (0,1)
You want y=mx+b; you need m and b
m=[-2-1]/[-2-0]=-3/-2=(3/2)
To find b, -2=(3/2)(-2)+b;
b-3=-2 ; b=1
y=(3/2)x+1
Cheers,
Stan H.